Demand schedules are given as follow
Derick: Qd=100-0.5p
Charles:Qd= 50-0.2p

What is the market demand schedule?
1
Expert's answer
2020-11-16T06:43:34-0500

To get the market demand function, we have the right to add the demand functions of two:

Qd=100-0.5p+50-0.2P=150-0.7P

let's draw three graphs





red Qd = 100-0,5p


blue Qd = 50-0.2p


green Qd=150-0.7P